OpenStudy (anonymous):

can you help me solve each equation system by the addition method? 4x+2y=6 4x-5y=-1 thank you. xx

Nnesha (nnesha):

hint: multiply 2nd equation by -1

Nnesha (nnesha):

either 1st or 2nd multiply one of equation by -1

OpenStudy (anonymous):

4x+2y=6 4x-5y=-1 Can we add the two and get rid of 4x? @Nnesha

OpenStudy (anonymous):

so then would it be 8x instead of just 4x?

OpenStudy (anonymous):

i think we can just add them and get off 4x

OpenStudy (anonymous):

that leaves us with 7y = 7

Nnesha (nnesha):

both 4x are positive

OpenStudy (anonymous):

the other one will become -4x though

OpenStudy (anonymous):

@satellite73 i am right or wrong?

OpenStudy (anonymous):

i mean like 4x + 2y = 6 4x - 5y = -1 so on adding it will be -4x _ 5y = 1

OpenStudy (anonymous):

-4x + 5y = 1 sorry

OpenStudy (anonymous):

so 4x + 2y = 6 -4x +5y = 1 that gives 7y = 7

Nnesha (nnesha):

@KenichiIvy yes that's why i said to multiply by -1

OpenStudy (anonymous):

So i am also correct @Nnesha ?

OpenStudy (anonymous):

7y = 7 so y = 7/7 y = 1

Nnesha (nnesha):

http://openstudy.com/code-of-conduct make sure u read this

OpenStudy (anonymous):

@Nnesha I am familiar with the rules.

OpenStudy (anonymous):

so would x be 5? im sorry, im aweful at algerbra.

OpenStudy (anonymous):

@alimcewen i suggest you to take a nice nap then try again

OpenStudy (anonymous):

and please remove the thought that you are awful at algebra

OpenStudy (anonymous):

@alimcewen

OpenStudy (raden_zaikaria):